OpenStudy (anonymous):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth?

OpenStudy (anonymous):

@srgntb0$$

OpenStudy (chunkymonkay):

4

OpenStudy (anonymous):

can you explain why?

OpenStudy (cwrw238):

you can make up an equation in x from this data let x be number of points for each multile question ; for Gretchen total points = 18x + 19 for Ezia total = 15x + 31 they both got same total points so 18x + 19 = 15x + 31 solve for x and you have the answer

OpenStudy (cwrw238):

18x - 15x = 31 - 19 3x = 12 x = 4

OpenStudy (chunkymonkay):

you will have to set this up into two equation. so the first equation would be 18x+19=y and the second one would be 15x+31=y 18x+19=15x+31 3x=12 x=4 which is your answer :)

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (chunkymonkay):

your welcome :)

OpenStudy (anonymous):

and thanks @cwrw238